867 is the product of two numbers one of which is 3 times greater than the other find the sum of the numbers ?
Answer by jorel555(1290)

Let n be your first number; then your 2nd number is 3n. So:
n(3n)=867
3n^2=867
n^2=289
n=17
3n=51
n+3n=68!!!!!!!!!!!!!!